# Language Models, Graph Searching, and Supervision Adulteration: When More Supervision is Less and How to Make More More

### v1

@@ 3.2.1 Ranking-into-the-Future (RITF)
We can go from LS to ranking-into-the-future by constructing rank targets and training with a rank-based loss, providing a structured loss over multiple tokens at each time step. As we are using future tokens, the structure is over the sequence and the sequential order is used for the rank-order.
@@
We incorporate another bias that ranks tokens in $R_{t}$ above all others, encoding the concept that the correct arm is more important than the others. [^6] This is not included in Eq. 2. See Appx. A.3. This creates very dense supervision with $M(M-1)/2$ intra- and $M^{2}(|V|-M)$ inner-arm pairs. Initial experimentation found this was crucial. This makes sense as, when the model fails to learn to predict $l_{t}$, it learns a uniform distribution over the set of leading nodes, meaning they have equal scores. This secondary bias creates supervision that $l_{t}$ is more important than the other leading nodes. Note that this is already implicitly done in any cross-entropy loss, including BOW and LS, as cross-entropy works by promoting the singular ground-truth while demoting all other nodes.
@@ 3.2.2 Results and Discussion
Exp. Desc. $D$ $M$ SR ABB
BoW 3 9 100% 100%
4 9 20% 60%
5 9 0% 0%
2 12 0% 20%
LS 3 7 100% 100%
3 9 0% 0%
RITF 4 9 100% 100%
5 9 60% 60%
3 12 100% 100%
4 12 60% 100%
2 15 60% 100%
Table 3: Alt. distro. results (full Tbl. 9 in Appx. A.3).
Tbl. 3 shows that RITF is superior to both BoW and LS. For LS we use a stepped monotonically decreasing weight (Frydenlund et al., 2022). We believe this does not work as it couples the inductive bias with loss scaling (so future tokens have tiny weights). We know the inductive bias is not at fault as it is the same as in RIFT. This shows that it is easier to specify rules than specific weights.
Future prediction works for multiple related reasons. First, the loss requires multi-edge lookups and so induces learning arm reconstruction. Second, it avoids adulteration by skipping adjacent inputs for all targets at $>i+1$. This is the same as masking, except instead of using noised input to cause the skip, it is implicitly defined as part of the loss. This is also fully masked thus inducing the desired backward alg. Third, by applying this at each time-step, the loss induces task decomposition across the sequence where learning $P_{B,\,i+1}$ is a sub-problem of learning $P_{B,\,i}$ (and is easier to boot since more conditioning input is provided).
@@ 3.3 Scratchpads (SP) to Increase Supervision
SPs predict an intermediate sequence before the target sequence, providing auxiliary input and target supervision (Nye et al., 2022). Both the reverse arm order and the arm-wise graph shuffle make the task trivial and so would be obvious SPs. These are problem-specific and do not prevent adulteration and, so, are not insightful. We present alternatives.
@@
We also use SPs which reconstruct the entire graph by ordering the arms (GR-SP). Full reconstruction would cause adulteration, so we just match leading- and target-node pairs. We order the arms by leading- or target-node value, again, introducing semantics (See Fig. 12 in Appx. A.5).
@@ 3.3.1 Results and Discussion
AR-SP results can be seen in Tbl. 10 in Appx. A.4. The reverse SP is trivially learnt as expected. The BoW SP starts failing to find the solution when $M=7$ and the ordered SP when $M=9$. While these make the PST learnable (on small scales), their performance is disappointing. However, their failures are informative. We report the sequential accuracy for $R_{t}$ and the SP separately inorder to evaluate where errors stem from. As Tbl. 10 shows, the models fail to correctly predict the SP and then fail to predict $R_{t}$ when conditioning on the incorrect SP. The BoW performance is informative as the obvious solution (to a human) would be to generate the BoW in the reverse order. Not learning to do this shows that the reverse solution is only trivial when it is provided with direct supervision (the same supervision that causes the CHC).
@@
GR-SP results can be seen in Tbl. 11 in Appx. A.5. These only learn to solve the task in 4/80 trials and this is exceedingly informative. We have four variants of the GR-SP as we can go from leading- to target-nodes or vice versa and then either sort by leading- or target-node values. Fig. 13 in Appx. A.5 plots the accuracy of each SP token across training. The models learn to correctly identify the needed sets of leading and target nodes. This is done by single-edge shortcuts; leading nodes by adjacency to $s$ and targets nodes by having no following edge. The models also correctly learn to sort either the leading- or target-nodes. This means that for the SP where we go from leading-to-targets, sorted by leading nodes the model can correctly identify the first leading-node but fails to connect it with its paired target. The same thing happens using targets-to-leading, sorted by target nodes.
@@
In both cases, the model knows which arm to reconstruct, and can condition on either the correct leading or target node, but still does not learn the actual reconstruction. Here all the model needs to do is deterministic path-following with no planning to choose the correct arm. This begs the question: if the solution is deterministic and does not require choices, is this actually a planning problem? These and the BoW SP results indicate that arm reconstruction is what makes the task hard – not planning. As we do not do the full graph reconstruction, we do not provide supervision for path-following or task decomposition. Thus these negative results are consistent with and indirectly support the theory that supervised task decomposition is necessary.
@@ 3.4.1 Results and Discussion
Exp. Desc. $D$ $M$ SR ABB
Query Subset 2 5 100% 100%
5 5 80% 80%
2 7 60% 100%
3 7 0% 0%
General Single Target 2 7 100% 100%
3 7 80% 100%
5 7 40% 40%
Table 4: Results for general query methods. Full Tbl. 12 in Appx. A.6, along with orginal-setting experiments.
Tbl. 4 shows that using a subset of $R_{t}$ makes the task learnable for small graphs. To verify that that is due to decreasing sensitivity, we tried sampling a general single node from $R_{t}$ to use as the target in $Q$. We find that not only does this make the task learnable, it performs better than using a subset. We expect this is because a subset introduces too much noise. However, if this does not work because of reduced sensitivity, why does it work? Because it induces task decomposition (illustrated in Fig. 4).
Hu et al. (2025a) performed this same experiment and found it did not learn the task. [^7] And they performed this experiment explicitly to determine if task decomposition makes the task learnable. To explain this contradiction, we experiment using the original task settings (using an offline dataset, $|V|=100$, and $Q$ after $G$). Here we find that the task is much harder to learn, with only 3/20 trials succeeding (Tbl. 12 in Appx. A.6). This implies that it would be easy to find only negative results, especially if seeded trials were not used. We argue further about issues with hyperparameters in Appx. A.6. This also highlights the importance of using an online dataset, and how other issues – which do not relate to planning – contribute to making the PST unlearnable in the original setting.
@@
3.5 From Path-Star to Tree-Star
We considered preventing the CHC by masking the target-side input. We also showed that we can induce task decomposition via general queries on the source-side. Here we partially prevent the CHC and induce task decomposition via another source-side modification; changing the graph topology.
We employ a trick to avoid this. By task definition, $t$ must be the last generated token. This precludes any subtree containing $t$ from being generated before the others. Then given a subtree containing $D^{\prime}$ child nodes (including $t$), the distribution over these children being valid continuations of a traversal is asymmetric in that $t$ is excluded but uniform over the remaining $D^{\prime}-1$ choices. However, this only works for subtrees containing $t$. This is achieved via LS over valid child nodes during training. Call these $D$-ary trees. See Fig. 5.
Figure 5: Red and green-dashed arrows form the desired pre-order traversal; red edges are included in $G$, while green are not, and hence are immune to the CHC.
Following this logic, we can design the tree to resolve any ordering ambiguity and avoid needing LS with a deterministic traversal by only allowing subtrees containing $t$ to have a max of 2 children and all others one child i.e. (structurally) lopsided binary trees. Call these split trees. See Fig. 6.
3.5.1 Results and Discussion
Tbl. 5 shows that training on split trees makes the task learnable. This can again be explained as inducing task decomposition. In Fig. 6, each subtask (marked as primes) is similar to a path-star graph with $D^{\prime}=2$ where a new start node, $s^{\prime}$, is any node with two children and the correct leading node, $l_{t}^{\prime}$, is the first node of the subtree not containing $t$. It is interesting that experiments where $D>2$ work since the induced subtask is restricted to $D^{\prime}=2$ and thus does not match the original task.
We find that the $D$-ary trees do not learn the task outside of a few cases, despite introducing a similar decomposition. We conjecture that smoothing over subtrees with $D$ choices creates a deficient subtask that mirrors the adulterated PST task since we are explicitly forcing the model to learn a uniform distribution over leading nodes and this matches the undesired learnt behaviour of models that fail the task i.e. we are teaching the model to do the very thing that we don’t want it to do.
Exp. Desc. $D$ $M$ SR ABB
Split Trees 2 5 100% 100%
5 5 100% 100%
2 7 60% 100%
3 7 100% 100%
5 7 60% 100%
2 9 80% 100%
5 9 0% 0%
Table 5: Split tree results (Tbl. 13 in Appx. A.7)
Training on trees and evaluating path-star graphs may look like an exotic solution, but we stress this is actually a generalization of graph topology and one that does not go far enough. We suspect that the best graph topology would be one that allows for perfect decomposition, where each subtask is the same as the original except for a change in the number of recurrent steps needed to solve the task i.e. the choice of graph topology will affect subtask homology. This conjecture would explain the necessity of ‘balanced’ graphs beyond preventing shortcuts (Saparov et al., 2025). [^8] Shortcuts are probably the symptom, not the illness.
Training on trees and then evaluating path-star graphs is also interesting since it is counterintuitive from the perspective of in-domain learning; we require training on trees to generalize to paths when direct training on paths fails. Ironically, the PST is defined as it is to rule out out-of-domain effects (Bachmann and Nagarajan, 2024). From the perspective of adulteration, paths are an overly informative graph structure. Also, training on paths to evaluate paths is more of a direct form of supervision than training on trees to evaluate paths.
@@ 3.6 Generalized Length Decomposition
Given the above, it should be obvious now that a direct way to induce task decomposition would be to supervise the training process by sampling different-sized graphs. Tbl. 14 in Appx. A.8 shows this makes the task learnable as expected. Training on various values of $D$ does not seem to help. Combining general length and target sampling improves performance over just doing the former. We expect these results would be better if given more training time, as this introduces a lot of noise.
@@ 4 Conclusion
We have taken the path-star task from being unlearnable to learnable with decoder-only models. We have shown how the original task is designed with adulterated supervision, explained why this makes it unlearnable due to the lack of decomposition supervision, and shown that preventing the CHC is not critical for learning the task given some decomposition supervision. We developed multiple methods to overcome this lack of supervision, with all retaining the next-token prediction paradigm with standard training via teacher-forcing. [^9] Except for the auxiliary future distributions.
We have empirically demonstrated that decomposition is critical for learning to search over graphs. This is strongly supported by the fact that the methods we have developed are all orthogonal to each other but can all be explained as inducing subtask decomposition in some form.
Our work serves as a bridge between Bachmann and Nagarajan (2024) and Saparov et al. (2025) by providing explanation for why the graph search task presented by the former is seemingly unlearnable, while the very similar graph search task presented by the latter is learnable; the latter provided decomposition supervision, while the former did not due to adulteration. Specifically, Saparov et al. (2025)’s task setup incorporated general queries, graph topologies, and lengths. [^10] They uniformly sampled $M$ and $D$ but not general queries when using path-star graphs. These still permit the CHC. While this is not critical for learning the task, preventing all shortcuts, say by masking, may be important for generalization.
Due to searching being recursively defined, subtask decomposition is inherent in the original task. Thus we can induce decomposition with the original supervision – provided we are careful not to adulterate it. This contrasts with other tasks that require introducing scratchpads with extra supervisory information. This is either done by modifying the task itself or learning a secondary subtask that is decompositional (Wies et al., 2023).
A few informative negative results suggest that the core difficulty of the task does not concern planning at all but rather graph reconstruction (and show that graph reconstruction is made more difficult for decoder-only models due to causal constraints via the causal-wise shuffling experiments). These also show that seemingly trivial solutions will not be found unless directly supervised.
If one was concerned about the implications that the empirical results of the path-star task had on the sufficiency of the next-token prediction paradigm for planning tasks, this work alleviates those concerns. If one is skeptical about such conjectures, this validates their beliefs with an explanation of why the PST in its original form is unlearnable. Indeed, our findings show that the task is fragile, where minor changes induce decomposition and make it leanrable, which indicates these issues will not apply to – or be as potent to – complex tasks.
@@ 5 Limitations
Scaling issues: The major limitation of our work is that each method fails to scale with either $D$ or $M$. We believe that using graph topologies that allow for stronger and more consistent decomposition where each subtask mirrors the main task will be key for scalability. Thus we think that path-star experimental setting is not a suitable environment to consider how the methods we have developed will scale to larger graphs. Instead, we believe using more general graphs would be a better environment and leave this to future work. However, Saparov et al. (2025) observed similar scaling issues using general graphs, which suggests that the issue does not stem from topology by itself (see Appx. B.2 for a comparison of our works). Still, this would let us avoid or isolate issues related to using path-star graphs when considering scale.
We conjecture that different issues may affect the scaling of $D$ and $M$. $M$ scales with the number of model layers (Frydenlund, 2024). It is unclear if learning the logarithmic algorithm is harder than the linear algorithm. As we use a model with 8 layers, and often find scaling past $M=9$ difficult, this may account for some scaling issues for $M$. We tied the vocabulary size to $D$ by setting $|V|=|G|$. This may account for some scaling issues for $D$ as this increases the sample space.
@@
The value of the path-star task: Given our findings, we argue that the PST is not a good task for evaluating the performance of general planning or search methods. We have introduced ranking-into-the-future but do not believe we have evaluated its full potential for planning tasks. Our discussion around RIFT is framed as a change in distribution from next-token prediction. While our goal of this work is to show that next-token prediction is sufficient for learning the task, we do not consider what is the best method. This also applies to the models developed by Yin et al. (2024); Hu et al. (2025a). [^11] They both perform additional planning experiments. We also argue that graph search is not a stand-in for general search. Given the finesse required to learn the PST and issues with scalability, we suggest that graph search may be hard due to issues that only apply to these experimental settings and caution against making claims based on these tasks being used as surrogates for more complex search tasks on natural data.
See below for positive uses of the task.
@@
Semantics as high-order graph structure: We explained how the path-star task is semanticless. This also applies to other related graph search tasks. While we do little experimentation in this direction, we conjecture that this plays an important role in learnability. In particular, the right semantics would probably make graph reconstruction significantly easier. This also relates to not attempting to use pretrained models, which would inherit natural language semantics from pertaining.
@@
Why is task decomposition necessary: We have empirically shown that task decomposition is necessary for learning the task. While it is intuitive why supervising decomposition will help learning, we do not explain why it is necessary. Thus characterizing the core underlying difficulty is still an open question. We expect solving this will be insightful for learnability theory of transformers. Thus we believe the original adulterated form of the path-star task is of scientific value and hope that research into it will continue.
While we did not find positive results trying to directly decrease the sensitivity by using multiple query nodes, our results are indecisive and may be explained by an increase in noise. Thus the sensitivity conjecture is still a possible explanation.
@@
Shortcuts: The PST is a great framework for exploring shortcut learning. Different variations result in different shortcuts, especially when using the SPs. One question we have is if shortcuts, once learnt, actively harm learning the desired task or if they are just benign symptoms of other issues.
@@
Alternative examples of adulteration: We describe adulteration as a generic issue but only consider the context of graph searching. However, we believe it is a useful term for identifying similar issues and showing they can be solved via similar methods. For example, Chang and Bisk (2025) considered trying to learn to count by providing a model with a contiguous sequence of numbers and training using next-token prediction. However, as they point out, this will be ineffective due to trivial bigram shortcuts. This is because they have presented the task in an adulterated form.

### v3

@@ 3.2.1 Ranking-into-the-Future (RITF)
We generalize from LS to ranking-into-the-future by constructing rank targets and training with a rank-based loss, providing a structured loss over multiple tokens at each time-step. As we are using future tokens, the structure is over the sequence, and the sequential order is used for the rank-order.
@@
We incorporate another bias that ranks tokens in $R_{t}$ above all others, encoding the concept that the correct arm is more important than the others. [^6] This is not included in Eq. 2. See Appx. B.3. This creates very dense supervision with $M(M-1)/2$ intra- and $M^{2}(|V|-M)$ inner-arm pairs. Initial experimentation found this was crucial. This makes sense as, when the model fails to learn to predict $l_{t}$, it learns a uniform distribution over the set of leading nodes, meaning they have equal scores. This secondary bias creates supervision that $l_{t}$ is more important than the other leading nodes. Note that this is already implicitly done in any cross-entropy loss, including BoW and LS, as cross-entropy works by promoting the singular ground-truth while demoting all other nodes.
@@ 3.2.2 Results and Discussion
Figure 6: Alt. distro. results (full Tbl. 4 in Appx. B.3).
Fig. 6 shows that RITF is superior to both BoW and LS. For LS we use a stepped monotonically decreasing weight (Frydenlund et al., 2022). We believe this does not work as it couples the inductive bias with loss scaling (so future tokens have tiny weights). We know the inductive bias is not at fault, as it is the same as in RITF. This shows that it is easier to specify rules than specific weights.
More importantly, future predictions make the task learnable for multiple related reasons. First, the multi-token loss requires multi-edge lookups and so induces decomposition. Second, it avoids adulteration by skipping adjacent inputs for all targets at $>i+1$. This is the same as masking, except instead of using noised input to cause the skip, it is implicitly defined as part of the loss. The first step is also fully masked, thus inducing the desired backward alg. for all tokens. Third, by applying this at each time-step, the loss induces novel dense decomposition across the sequence where learning $P_{B,\,i+1}$ is a sub-problem of learning $P_{B,\,i}$ (see Appx. A.2 for an expanded discussion).
@@ 3.3 Scratchpads (SP) to Increase Supervision
SPs predict an intermediate sequence before the target sequence, providing auxiliary input and target supervision (Nye et al., 2022). Both the reverse arm order and the arm-wise graph shuffle make the task trivial, and so would be obvious SPs. These are problem-specific and do not prevent adulteration and, so, are not insightful. We present alternatives.
@@
We also use SPs that reconstruct the entire graph by ordering the arms (GR-SP). Full reconstruction would cause adulteration, so we just match leading- and target-node pairs. We order the arms by leading- or target-node value, again, introducing semantics (see Fig. 21 in Appx. B.5).
@@ 3.3.1 Results and Discussion
AR-SP results can be seen in Tbl. 5 in Appx. B.4. We report the sequential accuracy for $R_{t}$ and the SP separately to isolate where any errors occur. The reverse SP is trivially learnt as expected. Some BoW trials are successful. Importantly, this can be explained by inducing subtask decomposition (Appx. A.3). While the BoW SP make the PST learnable (on small scales), its performance is disappointing, since an obvious solution (to a human) exists, which would be to generate the BoW in the reverse order, however, this solution is not found. As Tbl. 5 shows, the models fail to correctly predict the SP and then fails to predict $R_{t}$ when conditioning on the incorrect SP. This failure is informative as it shows the reverse solution is only trivial when it is provided with direct supervision (the same sequence supervision that causes the CHC).
@@
GR-SP results can be seen in Tbl. 6 in Appx. B.5. These only learn to solve the task in 4/80 trials and this is exceedingly informative. We have four variants of the GR-SP, as we can go from leading- to target-nodes or vice versa, and then either sort by leading- or target-node values. Fig. 22 in Appx. B.5 plots the accuracy of each SP token across training. The models learn to correctly identify the needed sets of leading and target nodes. This is done by single-edge shortcuts; leading nodes by adjacency to $s$ and targets nodes by counting degree. The models also correctly learn to sort either the leading- or target-nodes. This means that for the SP where we go from leading-to-targets, sorted by leading nodes, the model can correctly identify the first leading-node but fails to connect it with its paired target. The same thing happens using targets-to-leading, sorted by target nodes.
@@
In both cases, the model knows which arm to reconstruct, and can condition on either the correct leading or target node, but still does not learn the actual reconstruction! Here, all the model needs to do is deterministic path-following with no planning to choose the correct arm. This begs the question: if the solution is deterministic and does not require choices, is this actually a planning problem? These and the BoW SP results indicate that arm reconstruction is what makes the task hard – not planning. Thus, these negative results are consistent with and indirectly support the theory that supervised task decomposition is necessary, since the scratchpads the fail are those that do not induce decomposition.
3.4 From Path-Star to Tree-Star
Here, we partially prevent the CHC and induce task decomposition via the source-side modification of changing the graph topology from paths to trees.
We employ a trick to avoid this. By task definition, $t$ must be the last generated token. This precludes any subtree containing $t$ from being generated before the others. Then, given a subtree containing $D^{\prime}$ child nodes (including $t$), the distribution over these children being valid continuations of a traversal is asymmetric in that $t$ is excluded but uniform over the remaining $D^{\prime}-1$ choices. However, this only works for subtrees containing $t$. This is achieved via LS over valid child nodes during training. Call these $D$-ary trees. See Fig. 7.
Figure 7: Red and green-dashed arrows form the desired pre-order traversal; red edges are included in $G$, while green edges are not, and hence are immune to the CHC.
Following this logic, we can design the tree to resolve any ordering ambiguity and avoid needing LS with a deterministic traversal by only allowing subtrees containing $t$ to have a max of 2 children and all others one child, i.e., structurally-lopsided binary trees. Call these split trees. See Fig. 8.
3.4.1 Results and Discussion
Fig. 9 shows that training on split trees makes the task learnable. This can again be explained as inducing task decomposition. In Fig. 8, each subtask (marked as primes) is similar to a path-star graph with $D^{\prime}=2$ where a new start node, $s^{\prime}$, is any node with two children and the correct leading node, $l_{t}^{\prime}$, is the first node of the subtree not containing $t$. It is interesting that experiments where $D>2$ work since the induced subtask is restricted to $D^{\prime}=2$ and thus does not match the original task.
We find that the $D$-ary trees do not learn the task outside of a few cases, despite introducing a similar decomposition. We conjecture that smoothing over subtrees with $D$ choices creates a deficient subtask that mirrors the adulterated PST task since we are explicitly forcing the model to learn a uniform distribution over leading nodes and this matches the undesired learnt behaviour of models that fail the task, i.e., we are directly teaching the model to learn the exact behaviour we are trying to avoid.
Figure 9: Split tree results (Tbl. 8 in Appx. B.7)
Training on trees and evaluating path-star graphs may look like an exotic solution, but we stress this is actually a generalization of graph topology and one that does not go far enough. We suspect that the best graph topology would be one that allows for perfect decomposition, where each subtask exactly mirrors the core task except for a change in the number of recurrent steps needed to solve the task. That is, the choice of graph topology will affect subtask homology. This conjecture would explain the necessity of ‘balanced’ graphs beyond preventing shortcuts (Saparov et al., 2025). [^7] Shortcuts are probably the symptom, not the illness.
Training on trees and then evaluating path-star graphs is also interesting since it is counterintuitive from the perspective of in-domain learning; we require training on trees to generalize to paths when direct training on paths fails. Ironically, the PST is defined as it is to rule out out-of-domain effects (Bachmann and Nagarajan, 2024). From the perspective of adulteration, paths are an excessively informative graph structure. Also, training on paths to evaluate paths is more direct task supervision than training on trees to evaluate paths.
@@ 3.5.1 Results and Discussion
Figure 10: Results for general query methods. Full Tbl. 7 in Appx. B.6, along with original-setting experiments.
Fig. 10 shows that using a subset of $R_{t}$ makes the task learnable for small graphs. To verify that that is due to decreasing sensitivity, we tried sampling a single node as a general single target (GST) from $R_{t}$. We find that not only does this make the task learnable, it performs better than using a subset. We expect this is because a subset introduces too much noise. However, if this does not work because of reduced sensitivity, why does it work? Because it induces task decomposition (illustrated in Fig. 11).
Hu et al. (2025a) performed this same experiment and found it did not learn the task. [^8] And they performed this experiment explicitly to determine if subtask decomposition makes the task learnable. To explain this contradiction, we experiment using the original task settings (using an offline dataset, $|V|=100$, and $Q$ after $G$). Here we find that the task is much harder to learn, with only 3/20 trials succeeding (Tbl. 7 in Appx. B.6). This implies that it would be easy to find only negative results, especially if seeded trials were not used. We argue further about issues with hyperparameters in Appx. B.6. This highlights the importance of using an online dataset, and how other issues – not related to planning – contribute to the PST being unlearnable in the original experimental setting.
@@ 3.6 Generalized Length Decomposition
Given the above, it should be obvious now that a direct way to induce subtask decomposition would be to supervise the training process by sampling different-sized graphs. Tbl. 9 in Appx. B.8 shows this makes the task learnable as expected. Training on various values of $D$ does not seem to help. Combining general length and target sampling improves performance over just doing the former. We expect these results would be better if given more training time, as this introduces a lot of noise.
@@ 4 Conclusion
We have demonstrated that the path-star task, which was seemingly unlearnable, is actually learnable with decoder-only models. We have shown how the original task is designed with adulterated supervision, explained why this makes it unlearnable due to obscuring or absorbing decomposition supervision, and shown that preventing the CHC is not critical for learning the task given some decomposition supervision. We developed multiple methods to overcome this lack of supervision, with all retaining the next-token prediction paradigm with standard training via teacher-forcing. [^9] Except for the future distributions.
We have empirically demonstrated that decomposition is critical for learning to search over graphs. This is strongly supported by the fact that the methods we have developed are all orthogonal to each other but can all be explained as inducing subtask decomposition in some form. [^10] See Appx. A for a summary and comparison of methods.
A few informative negative results suggest that the core difficulty of the task does not concern planning at all but rather graph reconstruction and that seemingly trivial solutions will not be found unless directly supervised. We also show that graph reconstruction is made more difficult for decoder-only models due to causal constraints via the causal-wise shuffling experiments.
Our work serves as a bridge between Bachmann and Nagarajan (2024) and Saparov et al. (2025) by providing explanation for why the graph search task presented by the former is seemingly unlearnable, while the very similar graph search task presented by the latter is learnable; the latter provided decomposition supervision (specifically they incorporated general graph topologies and lengths), while the former did not due to adulteration. Note, these still permit the CHC. While this is not critical for learning the task, preventing all shortcuts, say by masking, may be important for generalization.
As searching is recursively defined, decomposition is inherent in the original task. Thus, we can induce it using the original supervision – provided we are careful not to adulterate it. This contrasts with other tasks that require introducing scratchpads or extra supervisory information – done by modifying the task itself or learning a secondary subtask that is decompositional (Wies et al., 2023).
If one were concerned about the implications that the empirical results of the path-star task had on the sufficiency of the next-token prediction paradigm for planning tasks, this work alleviates those concerns. If one is skeptical about such conjectures, this validates their beliefs with an explanation of why the PST in its original form is unlearnable. Our findings show that the task is fragile, where minor changes induce decomposition and make it leanrable, which indicates these issues will not apply to – or be nearly as potent to – complex tasks as conjectured.
@@ 5 Limitations
Scaling issues: The major limitation of our work is that each method fails to scale with either $D$ or $M$. We believe that using graph topologies that allow for stronger and more consistent decomposition, where each subtask mirrors the main task, will be key for scalability. Thus, we think that path-star experimental setting is not a suitable environment to consider how the methods we have developed will scale to larger graphs. Instead, we believe using more general graphs would be a better environment and leave this to future work. However, Saparov et al. (2025) observed similar scaling issues using more general graphs, which suggests that the issue does not stem from topology by itself (see Appx. C.2 for a comparison of our works). Still, this would let us avoid or isolate issues related to using topology when considering scale.
We conjecture that different issues may affect the scaling of $D$ and $M$. $M$ scales with the number of model layers (Frydenlund, 2024). It is unclear if learning the log alg. is harder than the linear alg.
@@
The value of the path-star task: Given our findings, we argue that the PST is not suitable for evaluating the performance of novel methods or architectures for planning (Yin et al., 2024; Hu et al., 2025a, among others). [^11] They also both perform additional non-PST experiments. We also argue that graph search is not a stand-in for general search. Given the finesse required to learn the PST and issues with scalability, we suggest that graph search may be hard due to issues that only apply to these experimental settings and caution against making claims based on these tasks being used as surrogates for other search tasks on natural data (see below for positive uses of the task). While our goal of this work is to show that next-token prediction is sufficient for learning the task, we do not consider what method is best. Given the above, we have introduced the novel ranking-into-the-future method, but have not evaluated its potential for planning on a suitable task. We frame RITF as a change in distribution from next-token prediction. This is important to the discussion of the paradigm in this work, but may not be important for other uses.
@@
Semantics as high-order graph structure: The path-star task is semanticless. This often applies to other minimal graph search tasks. While we do little experimentation in this direction, we conjecture that this plays an important role in learnability, where the right semantics would probably make graph reconstruction significantly easier. This relates to not using pretrained models, which would inherit natural language semantics.
@@
Why is task decomposition necessary: We have empirically shown that task decomposition is necessary for learning the task. While it is intuitive why supervising decomposition will help learning, we do not explain why it is necessary. Thus, characterizing the core underlying difficulty is still an open question. We expect solving this will be insightful for learnability theory. Thus we believe the original adulterated form of the path-star task is of scientific value and hope that research into it will continue. We did not find positive results trying to decrease sensitivity with multiple query nodes, but our results are inconclusive and may be due to an increase in noise. Thus, the sensitivity conjecture is still a possible explanation.
@@
Shortcuts: The PST is a great framework for exploring shortcut learning. Different variations result in different shortcuts, especially when using the SPs. One question we have is if shortcuts, once learnt, actively harm learning the desired task or if they are just benign symptoms of other issues?
@@
Alternative examples of adulteration: We describe adulteration as a generic issue but only consider the context of graph searching. We believe it is a useful term for identifying similar issues which can be solved via similar methods. For example, Chang and Bisk (2025) considered trying to learn to count by providing a model with a contiguous sequence of numbers and training using next-token prediction. As they point out, this will be ineffective due to trivial bigram shortcuts. This is because they have presented the task in an adulterated form.